# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v1

@@ 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. The following three kinds of $\mathcal{A}$-valued kernels are often considered in existing studies [14, 13, 1, 9, 4].
@@
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the existing kernels. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
@@
With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, let
@@
$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (1)
@@
$F_{n}^{q,P}(t)=1/n\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}$ (2)
for $t\in\mathbb{T}^{q}$ is the Fejér kernel on $\mathbb{T}^{q}$ with respect to the polyhedron $P$. Here, for $j\in\mathbb{Z}$, $jP=\{jz\,\mid\,z\in P\}$. The generalization of the Fejér kernel on $\mathbb{T}$, which is defined in Subsection 2.2, to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated [26, 25]. Note that $F_{n}^{q,P}(t)$ is a real-valued function since for $r\in P$, we have $-r\in P$. Thus, if a kernel in Example 2 is real-valued, then the corresponding kernel in Definition 6 is also real-valued. Eq. (1) implies that the value of $S_{n}\big{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y_{j})\big{)}$ at $z$ is determined not only by $x_{1}(z),\ldots,x_{q}(z),y_{1}(z),\ldots,y_{q}(z)$, but also by $x_{1}(t_{1}),\ldots,x_{q}(t_{q}),y_{1}(t_{q+1}),\ldots,y_{q}(t_{2q})$ with different values $t_{1},\ldots,t_{2q}$ from $z$.
@@ 3.1 Convergence and interactions
We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2, respectively. In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of existing kernels. In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$. In the following, all the proofs of theoretical statements are documented in Appendix A.
@@
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 8 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (1). Lemma 9 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini and Travaglini [25]. We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (1), and apply Lemmas 8 and 9 to show Theorem 7. See Appendix A for more details.
@@ 3.2 Positive definiteness
To define an RKHM introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product. Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
@@
The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
As for $k_{n}^{\operatorname{prod},q}(x,y)$, $x$ and $y$ depend on $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products. Thus, we cannot show its positive definiteness. However, we can modify it to become positive definite as follows.
@@
Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
@@
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, the bound in Lemma 13 may not be tight, and even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated [27, 28]. Indeed, in practical computations in Subsection 5.1, the algorithms can work even if we set $\beta_{n}<n^{q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@ 4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. To derive the generalization bound, we derive an upper bound of the Rademacher complexity. Let $\Omega$ be a probability space with the probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be input and output samples from a distributions of $\mathcal{A}_{0}^{d}$-valued and $\mathcal{A}_{1}$-valued random variables $X$ and $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d. Rademacher variables, which take their values on $\{-1,1\}$. For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$. Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$. An $\mathcal{A}$-valued version of the Rademacher complexity is also defined by Hashimoto et al. [14] as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$. Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3.
@@
Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
The first term of the bound (3) is the empirical loss, and the second term of the bound represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$ associated with $k_{n}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$ in $\mathcal{A}$ is larger, $\mathcal{M}_{k_{n}}$ can represent more functions. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, according to Theorem 15, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger. An advantage of introducing the parameter $n$ and constructing the new kernels based on the existing kernels in Example 2 is that we can control the empirical loss and the complexity through $n$. By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$ small.
To show Theorem 15, we use the $\mathcal{A}$-valued Rademacher complexity. Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity.
Proposition 16
@@ 5 Application to Kernel Ridge Regression
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$. Consider the typical minimization problem for regression, which is also considered in [14]:
@@
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. On the other hand, that for $k^{\operatorname{poly},q}$, $\hat{k}^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ is $O(qmN^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use a vvRKHS. For vvRKHSs, to obtain $m$ values as an output, we generally apply a $\mathbb{C}^{m\times m}$-valued kernel. In this case, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$. We can reduce the cost with respect to $N$, e.g., by Nyström method [29]. Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
5.1 Numerical results

### v3

@@ 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. To obtain vector- or functional-valued outputs, applying vvRKHSs has been investigated. However, as we stated in Section 1, to go beyond the separable kernel and reach higher performance, the computational cost scales as $O(m^{3}N^{3})$ for $m$-dimensional outputs with $N$ samples. To resolve this issue, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs. As we will discuss in the last part of Section 5, the computational cost depends on $mN^{3}$ if we use an $\mathcal{A}$-valued kernel. Hashimoto et al. (2021) proposed to use the $\mathcal{A}$-valued kernel $k$ where $k(\cdot,\cdot)(z)$ is a complex-valued positive definite kernel for all $z\in\mathbb{T}$. Combining it with typical existing kernels (Hashimoto et al., 2023a; Schölkopf & Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012), we obtain the following three kinds of $\mathcal{A}$-valued kernels.
@@
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the typical kernels in Example 2. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
@@
With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$, let
@@
Remark 7
Although the inputs of the kernels in Example 2 and Definition 6 are functions, we can also deal with vector inputs. For vector inputs, we can regard them as values of functions, and approximate the integral by the discrete sum.
@@
$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\!\!\!\!\!\!\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (2)
@@ 3.1 Convergence and interactions
We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2 as $n$ goes to infinity, respectively. In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of typical kernels. In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$. In the following, all the proofs of theoretical statements are documented in Appendix A.
@@
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 9 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (2). Lemma 10 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini & Travaglini (1997). We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (2), and apply Lemmas 9 and 10 to show Theorem 8. See Appendix A for more details.
@@ 3.2 Positive definiteness
To construct an RKHM, which is introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product. Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
@@
The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
As for $k_{n}^{\operatorname{prod},q}(x,y)$, it depends on $x$ and $y$ through $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products. Thus, we cannot show its positive definiteness. However, we can modify the kernel to become positive definite as follows.
@@
Let $\beta_{n}\geq-\min_{z\in\mathbb{T}^{q}}F_{n}^{2q,P}(z)$. Then, $\hat{k}_{n}^{\operatorname{prod},q}$ defined below is positive definite.
@@
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated (Ong et al., 2004; Canu et al., 2005). Indeed, in practical computations in Subsection 7, the algorithms can work even if we set $\beta_{n}<n^{2q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@ 4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity associated with the proposed kernels. According to Lemma 4.2 in Mohri et al. (2012), the generalization bound is described by the Rademacher complexity. We apply this lemma and derive a generalization bound.
Let $\Omega$ be a probability space with a probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\mathrm{E}$ be the Bochner integral on $\Omega$ with respect to $\mu$. Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3.
@@
Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
The first term of the right-hand side of (3) is the empirical loss, and the second term represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ in $\mathcal{A}$ is larger, the outputs of ${k_{n}}$ can represent more functions, which enables us to describe the dependency of outputs on inputs more. Indeed, if $n=1$, then the output of $k_{n}$ is always a constant function, where the situation is the same as the existing separable kernel discussed in Section 1. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, according to Theorem 16, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger. An advantage of introducing the parameter $n$ to construct the new kernels based on the typical kernels in Example 2 is that we can control the empirical loss and the complexity through $n$. By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$, the left-hand side of Eq. (3), small.
@@ 5 Application to Kernel Ridge Regression
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$. Consider the typical minimization problem for regression, which is also considered by Hashimoto et al. (2023a):
@@
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use vvRKHSs. For vvRKHSs, to obtain $m$ values as an output, we apply a $\mathbb{C}^{m\times m}$-valued kernel. If the kernel is nonseparable, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$. We can reduce the cost with respect to $N$, e.g., by Nyström method (Drineas & Mahoney, 2005). Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
6 Deep model with the spectral truncation kernels
As we discussed in Section 4, there is a tradeoff between the representation power and the model complexity. In fact, there are two directions of the representation power: one is about the dependency of the outputs on the inputs discussed in Section 4, and the other is about the interactions along the data function domain. One shortcoming of the spectral truncation kernel is that to obtain a model with a high representation power in the sense of the dependency of the outputs on the inputs, the truncation parameter $n$ have to be large, which results in making the interactions along the data function domain weak. To gain the two types of representation powers at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n}^{1},\ldots,k_{n}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (6)
We consider the case where $c_{i}^{j}$ in Eq. (6) is represented as a finite sum of functions in $\mathcal{A}$. In the following case, the representation power of this model grows exponentially.
Proposition 17
Let $k^{j}_{n}(x,x_{i})=\sum_{l=-(n-1)}^{n-1}\tilde{d}_{i,l}^{j}\mathrm{e}^{\mathrm{i}lz}$ with $\tilde{d}_{i,l}^{j}\in\mathbb{C}$. Assume $c_{i}^{j}$ is parameterized as $c_{i}^{j}(z)=\sum_{l=-(n-1)}^{n-1}d_{i,l}^{j}e_{\tau_{j},l}$ with $d_{i,l}^{j}\in\mathbb{C}$ and $\tau_{j}\in\mathbb{R}$. Here, $e_{\tau_{j},l}\in C(\mathbb{T})$ satisfies $e_{\tau_{j},l}(z)=\mathrm{e}^{\mathrm{i}\tau_{j}lz}$ for $z\in S$ for a subset $S$ of $\mathbb{T}$. Then, for $z\in S$, $f$ in Eq. (6) is represented as
$\sum_{l_{1},\ldots,l_{L}=-(n-1)}^{n-1}\sum_{m_{1},\ldots,m_{L}=-(n-1)}^{n-1}\sum_{i=1}^{N}\tilde{d}_{i,l_{1}}^{1}\cdots\tilde{d}_{i,l_{L}}^{L}{d}_{i,l_{1}}^{1}\cdots{d}_{i,l_{L}}^{L}\mathrm{e}^{\mathrm{i}(l_{1}+\cdots+l_{L}+\tau_{1}m_{1}+\cdots\tau_{L}m_{L})z}.$
Let $V_{n,L}=\operatorname{Span}\{z\mapsto\mathrm{e}^{\mathrm{i}(l_{1}+\cdots+l_{L})z}e_{\tau_{1},m_{1}}(z)\cdots e_{\tau_{L},m_{L}}(z)\,\mid\,l_{1},\ldots,l_{L},m_{1},\ldots,m_{L}\in\{-(n-1),\ldots,n-1\}\}$. If $\{\tau_{1},\ldots,\tau_{L}\}$ is linearly independent over $\mathbb{Z}$, then $V_{n,L}$ is described by $2n+1+(2n+1)^{L}$ oscillating functions. In this sense, the representation power of the model $f$ grows exponentially with respect to the number of layers $L$.
Remark 18
The exponential growth of the representation power is also observed for neural networks (Hanin & Rolnick, 2019). Deep learning with kernels has also been proposed (Laforgue et al., 2019; Hasimoto et al., 2023). Whereas the growth is obtained by the composition for these existing frameworks, it is obtained by the product for our case. Unlike the case of the existing frameworks, the representation power regarding $\mathcal{A}$ does not become high by the composition of functions.
Remark 19
The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k^{j}$. Since the Rademacher complexity bound depends on the norm of the kernel, it can also grow exponentially with respect to $L$. Thus, for the deep setting, in the same manner as Theorem 16, we still have the tradeoff between the representation power and the model complexity. In this case, it is controlled by both $n$ and $L$. The exponential growth of the Rademacher complexity bound (thus, the generalization bound) with respect to the number of layers is also observed for neural networks (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018). Investigating how we can take advantage of the deep setting in the sense of the generalization is future work.
7 Numerical results